Erin is on a train traveling at 70 mph. She spots a train traveling at 60 mph and records it takes 6 seconds for the 2nd train to pass. How long (in feet) is the second train if the trains are going in the same direction?
Answer by mananth(16946)

1 mph = 5280/3600
=22/15 feet/s
difference in speed = 10 mph
10 mph = 10 *22/15 feet/s
44/3 feet/s
d= rate * time
d=44/3 * 6
= 88 feet
length of train = 88 feet